### v2

# Distance measurements from the internal dynamics of globular clusters: Application to the Sombrero galaxy (M104)

Abstract
Globular clusters (GCs) are dense star clusters found in all massive galaxies. Recent work has established that they follow a tight relation between their internal stellar velocity dispersion and luminosity, enabling accurate distance measurements. In this work, we aim to apply this GC velocity dispersion (GCVD) distance method to measure the distance to M104 (NGC 4594, the Sombrero galaxy). We have measured internal stellar velocity dispersions for 85 globular clusters (GCs) and one ultra-compact dwarf galaxy around M104 using high-resolution multi-object integrated-light spectroscopy with FLAMES/GIRAFFE on the Very Large Telescope. The measured velocity dispersions range from km s-1, with a mean uncertainty of km s-1. For a subset of 77 GCs with -band magnitudes and reliable velocity dispersion measurements above km s-1, we constructed the - relation to measure the distance to M104, finding (stat.) (sys.) Mpc. The GCs follow the Milky Way and M31 relation closely, with the exception of the luminous ultra-compact dwarf SUCD1, which is nearly one magnitude brighter than the mean relation. 29 GCs in the sample have sizes determined from Hubble Space Telescope imaging which allowed us to determine their masses and -band dynamical mass-to-light ratios (M/LV). We find a mean for the luminous ( mag) M104 GCs, which is higher than the Milky Way GCs, but is reminiscent of the brightest GCs in Centaurus A. With the exception of SUCD1, the GCs of M104 follow the GCVD relation irrespective of their mass-to-light ratio.
keywords:
galaxies: distances – galaxies: star clusters – galaxies: M1041 Introduction
Globular clusters (GCs) are ancient star clusters found in all massive galaxies (Brodie and Strader, 2006; Beasley, 2020). With stellar masses and compact sizes ( pc), they are dense enough to have survived for billions of years. Consequently, as some of the oldest stellar structures in the Universe, GCs have long been employed as ancient fossils of galaxy formation and assembly.
In the Milky Way, spatially resolved studies have placed constraints on the chemical and kinematic properties of individual stars within GCs. For decades it has been known that GCs follow a tight relation between their magnitudes (or masses) and internal velocity dispersions (Pryor and Meylan, 1993). Similar to the Faber-Jackson relation for elliptical galaxies (Faber and Jackson, 1976), this relation is a projection of the fundamental plane that arises from the Virial Theorem (Djorgovski, 1995; Strader et al., 2009): In a self-gravitating system, the total mass scales with , where refers to the half-light radius. As old stellar systems with a narrow size distribution (pc Masters et al. 2010) and a narrow range of mass-to-light ratios (1 - 3; McLaughlin 2000; Baumgardt and Hilker 2018), the GC relation shows remarkably small scatter. As already discussed by Paturel and Garnier (1992), Djorgovski (1995), and Strader et al. (2009), this tight relation between GC magnitudes and velocity dispersion can be employed as an accurate distance indicator.
Recently, Beasley et al. (2024) revisited the GC velocity dispersion (GCVD) distance relation by collecting -band magnitudes and velocity dispersions for GCs in the Milky Way and M31 (Baumgardt and Hilker, 2018; Strader et al., 2011) in order to recalibrate the relation using geometric distances derived from Gaia parallaxes. By then applying the GCVD relation to Centaurus A using 57 GCs with velocity dispersion measurements from Dumont et al. (2022), Beasley et al. (2024) demonstrated that this method can yield distances with uncertainties as low as . Even more recently, the GCVD method was applied to derive a distance to the ultra diffuse galaxy NGC 1052-DF2, achieving an uncertainty of based on 5 GCs (Beasley et al., 2025).
As of now, application of the GCVD distance method is limited by the availability of velocity dispersion measurements. Outside the Local Group, dispersion measurements are rare as they rely on moderately high spectral resolution () to measure dispersions down to km s-1, a typical value for a GC with a mass of (the approximate mean mass of GCs in the Milky Way, Baumgardt et al. 2020). In an ongoing effort to collect larger samples of GC velocity dispersions, we present here velocity dispersions of 85 GCs and one ultra-compact dwarf galaxy (UCD) around M104 (NGC 4594, the Sombrero galaxy), based on high-resolution spectroscopy with the FLAMES GIRAFFE multi-object spectrograph at the Very Large Telescope (VLT). In Fahrion et al. (2025), we presented an extensive catalogue of GC velocities using the same dataset combined with spectra from the OSIRIS instrument at the Gran Telescopio Canarias (GTC) and analysis of archival VLT MUSE data. In this paper, our focus is on deriving a GCVD distance to M104. Additionally, for a subsample of 29 GCs with available size measurements, we derive dynamical masses and mass-to-light ratios, allowing us to test the effect of sizes and mass-to-light ratios on the plane.
M104 is a particularly interesting target for the GCVD method. As a peculiar galaxy with its prominent, almost edge-on disc, M104 has been studied extensively using space- and ground-based facilities (e.g. Emsellem 1995; Harris et al. 2010; Jardel et al. 2011; Gadotti and Sánchez-Janssen 2012; Cohen et al. 2020). Also, its rich GC system has been the focus of many studies (e.g. Wakamatsu 1977; Harris et al. 1984; Bridges and Hanes 1992; Bridges et al. 1997; Forbes et al. 1997; Larsen et al. 2002; Rhode and Zepf 2004; Bridges et al. 2007; Alves-Brito et al. 2011; Dowell et al. 2014; Kang et al. 2022). The total number of GCs is estimated to lie between 1500 and 2000 (Rhode and Zepf, 2004; Kang et al., 2022), and combining literature samples with newly collected data, we collected line-of-sight velocities for 499 of those in Fahrion et al. (2025), corresponding to approximately 25% of the total GC system. With such a rich GC system, M104 is therefore an ideal target to apply the GCVD method and test it against other distance measurements.
Distance measurements to M104 from the literature range between 7 and 11 Mpc111From the NASA/IPAC Extragalactic Database. Additional estimates using the Tully-Fisher relation give generally larger distances ( 11 - 22 Mpc; Bottinelli et al. 1984; Tully and Fisher 1988; Sorce et al. 2014), but it is questionable if this relation can be applied to M104 with its peculiar morphology and gas distribution (see e.g. McQuinn et al. 2016). and encompass a variety of different distance indicators. Using the tip of the red giant branch (TRGB) in deep Hubble Space Telescope (HST) observations, McQuinn et al. (2016) derived a distance of (sys.) Mpc, which is the most recent precise distance measurement to M104. Older distance measurements using surface brightness fluctuations (SBF) based on infrared HST observations, Jensen et al. (2003) reported a distance of 9.08 Mpc, similar to other SBF measurements (Ciardullo et al., 1993; Ferrarese et al., 2000; Tonry et al., 2001) that found distances between 8 and 10 Mpc. Similar values were also reported based on the luminosity functions of GCs (GCLF; Bridges and Hanes 1992; Spitler et al. 2006) and planetary nebulae (PNLF; Ciardullo et al. 1993; Ford et al. 1996).
The aim of this paper is to construct the GCVD relation for M104 GCs for a precise distance measurement. The paper is structured as follows: In Sect. 2, we describe the FLAMES high-resolution data. Sect. 3 describes our methods of measuring dispersions, the available photometry and our aperture correction. Sect. 4 describes the distribution of M104 GCs on the - plane. The distance measurement is detailed in Sect. 5 and compared to literature distances in Sect. 6. We summarise and conclude in Sect. 7.
2 Data
[…]
Full details of the observations and data reduction are given in Fahrion et al. (2025). To summarise, we used the FLAMES GIRAFFE HR21 (875nm) grating, which provides coverage of the calcium triplet region from 8483 to 8900 Å with resolution of and sampling of 0.05 Å pix-1. Each pointing was covered in ten exposures of 1365 seconds each, distributed over five observing blocks. Targets were selected from the catalogues of Spitler et al. (2006), Harris et al. (2010) and Kang et al. (2022).
The data were reduced using standard procedures using the FLAMES GIRAFFE esorex pipeline and custom scripts, with the variation that the sky-subtraction was performed using pPXF (Cappellari and Emsellem, 2004; Cappellari, 2017, 2023) to fit each spectrum with a stellar model from the PHOENIX library (Husser et al., 2013) and sky templates as sky input. After sky subtraction, spectra from individual exposures were corrected for barycentric velocities, resampled, and combined using a median weighted by the signal-to-noise (S/N) ratio of each spectrum.
We also kept a version of spectra with sky lines for masking purposes (see below). Additionally, we used those to test the final spectral resolution of our spectra by fitting prominent sky lines with Gaussian curves. We found a median full width at half maximum (FWHM) of the sky lines of 0.54 0.04 Å. and did not observe any variation in the spectral resolution with wavelength.
3 Methods
In the following, we describe our methods to derive velocity dispersions from the FLAMES spectra. Additionally, we describe the available photometry and how mass-to-light ratios were derived. Sect. 5 then details the distance measurements.
3.1 Spectral fitting
We fit the FLAMES GC spectra with pPXF (Cappellari and Emsellem, 2004; Cappellari, 2017, 2023). pPXF fits the spectrum with a library of suitable model or empirical spectra and returns the parameters of the line-of-sight velocity distribution (LOSVD) such as radial velocity and velocity dispersion. We chose additive polynomials of degree 4 and multiplicative polynomials of degree 1 to account for variations in the continuum and flux calibrations, but as described in Beasley et al. (2025), the results are not very sensitive to the choice of degree of the polynomials. Because sky residuals remained in the reduced data, we carefully masked those lines using a custom mask for every spectrum. We began by normalising the spectrum without sky subtraction in a continuum wavelength range to the continuum level of the sky-subtracted spectrum. Then, we masked out all wavelength pixels in which the normalised sky spectrum exceeds 1.2 times the continuum level. We fitted all spectra in a wavelength range between 8450 and 8850 Å except for four GCs that had strong sky residuals at redder wavelengths. Here, we restricted the upper wavelength range to 8750 Å due to strong sky residuals at longer wavelengths.
We used the high-resolution stellar population models from Coelho (2014) to fit the FLAMES spectra. In a first step, we selected red (sub) giant stars with effective temperature between 4500 and 5000 K, surface gravity between log(g)= 2.0 and 4.0, and metallicities between 1.5 and 0 dex. Then, the selected model spectra were brought to a spectral resolution of 0.54 Å to match the FLAMES spectra. Figure 1 shows eight GC spectra and their fits as examples, chosen to cover a range of -band magnitudes and signal-to-noise ratios.
As described in Fahrion et al. (2025), we performed a Monte-Carlo approach to obtain random uncertainties. For this, 100 representations of the input spectrum were created by drawing values from the residual of the initial fit and adding it onto the initial best fit spectrum. These 100 spectra where then fitted with pPXF for bootstrapping. We used the corresponding histograms of line-of-sight velocities and dispersions to derive the best-fit values and their uncertainties as the 16th, 50th, and 84th percentiles of the distributions. We show the resulting uncertainties on the velocity dispersion as a function of spectral signal-to-noise (S/N) and GC -band magnitude (from Kang et al. 2022) in Fig. 2. From the full sample of 93 GCs, 85 have velocity dispersion uncertainties below 5 km s-1 and dispersions km s-1. For the distance measurements, we considered this sample of 85 GCs with reliable velocity dispersion measurements ( km s-1, km s-1).
3.2 Aperture Corrections
[…]
We give the full sample of velocity dispersion measurements in Appendix B. The values range between 4 and 30 km s-1 with mean uncertainties of 2.5 km s-1. For SUCD1, we measured an aperture-corrected dispersion of km s-1, in agreement with the measurement by Hau et al. (2009) of km s-1 based on a lower resolution Keck/DEIMOS spectrum.
3.3 Available photometry
To directly apply the GCVD approach using the relation calibrated on Local Group GCs, -band magnitudes are needed. For our sample of 85 GCs (not including SUCD1) with velocity dispersion uncertainties km s-1 and velocity dispersions km s-1, 29 have high quality -band magnitudes available from imaging with the Advanced Camera for Surveys (ACS) onboard HST data (Spitler et al., 2006; Harris et al., 2010).
Moreover, 68 GCs have -band magnitudes available from Rhode and Zepf (2004) based on imaging with the Mosaic detector on the Kitt Peak 4 metre telescope. The ones with -band magnitudes available from both studies show good agreement within 0.1 mag. In total, -band magnitudes are available for 77 of 85 suitable GCs, in a magnitude range of mag. For SUCD1, Hau et al. (2009) reported a -band magnitude of 17.46 mag based on the same HST/ACS observations as used by Spitler et al. (2006) and Harris et al. (2010). In addition, all GCs have -band magnitudes from Kang et al. (2022) available, based on CFHT/MegaCam imaging.
Figure 3 shows the relation between velocity dispersion and apparent magnitudes in different filters. In all cases, a strong relationship is found due to the underlying relation between mass and dispersion. As the GCVD relation is currently only calibrated on -band magnitudes (shown in the top left panel), we show here the other filters only for visualisation. In Sect. 5 we also explore distance measurements based conversions from and to -band magnitudes.
4 The Sombrero GCs on the - plane
Before deriving a GCVD distance, we explore the properties of the GC sample in the - plane to test underlying dependencies with GC parameters. In Fig. 4, we colour-coded the GCs according to different parameters to test if the scatter correlates with GC parameters. For the full sample, we collected -band magnitudes from the HST study by Harris et al. (2010) and ground-based measurements from Rhode and Zepf (2004).
4.1 Comparison to other GC samples
The top left panel of Fig. 4 compares the band magnitudes and velocity dispersions of the M104 GCs with that of the Milky Way and M31 (white circles), which were used to calibrate the GCVD relation (Beasley et al., 2024), assuming a distance of 9 Mpc. Additionally, GCs in Centaurus A are shown with orange triangles (Dumont et al., 2022). As this panel shows, the M104 GCs cover a very similar range in magnitudes and dispersions as the Cen A GCs. The scatter in each direction appears similar to the calibration sample. SUCD1 is about a magnitude brighter than even the brightest GC in the calibration sample or Cen A.
4.2 Metallicities, colours and velocities
In the second panel on the top left of Fig. 4, the GCs are colour-coded according to their metallicity. The metallicity values are drawn from the full M104 sample described in Fahrion et al. (2025). Where available, we used metallicities from optical VLT/MUSE or GTC/OSIRIS spectroscopy, and for the remaining GCs the metallicities were derived directly from full spectrum fitting to the FLAMES spectra. While the GCs with measured dispersions cover a broad range of metallicities, there seems to be no clear correlation with the scatter in the magnitude-dispersion plane. Splitting the MW and M31 GCs into metal-poor and metal-rich GCs, Beasley et al. (2024) found no difference in the GCVD relation’s slope or offset, in agreement with our finding for M104.
Similarly, there seems to be no clear correlation between the colour and the scatter on the GCVD relation, as the second panel of Fig. 4 shows. Also, the scatter does not seem to correlate with the line-of-sight velocity of the GCs.
4.3 Projected distances
The bottom left panel in Fig. 4 shows the GCs colour-coded by their projected distance to the centre of M104. The projected distances range from 1 to 20, and no obvious relation with the scatter in the - plane is apparent. We note that the brightest GCs in our sample are also relatively close (a few arcminutes) to the centre. This could be a selection effect or could reflect mass segregation of the GC sample since more massive GCs tend to experience stronger dynamical friction effects (e.g. Tremaine 1976; Turner et al. 2012; Lotz et al. 2001). In the Appendix A we show the magnitudes and dispersions as a function of galactocentric distance.
4.4 Globular cluster sizes
As the GCVD relation is a consequence of the Virial Theorem, based on the relation between total mass and the combination of velocity dispersion squared and size (), the size influences the position of a source on the relation, with the effect expected to scale linearly. This is seen in the second panel in the bottom row of Fig. 4, where smaller-sized GCs lie to the right of the GCVD relation. Also, the observed offset of SUCD1 could be explained by the size difference as it is a particularly large object with a reported half-light radius of pc (Hau et al., 2009), significantly larger than typical GCs (mean pc, Harris et al. 2010; Masters et al. 2010).
Therefore, to test if the GCVD relation is applicable to M104, we compare the sizes of its GCs with the Milky Way (from Baumgardt and Hilker 2018) and M31 GCs (from Barmby et al. 2007; Peacock et al. 2010; Strader et al. 2011) that were used for the original calibration of the relation (Beasley et al., 2024). Figure 5 shows this comparison, based on the size measurements of 652 M104 GCs from Harris et al. (2010) which were based on HST ACS imaging. We find that the size distribution of M104 GCs aligns well with the sizes of the calibration MW and M31 sample.
However, we note that the small subsample of 29 GCs for which we have both sizes and dispersion measurements (shown in pink in Fig. 5) contains preferentially larger GCs. As these GCs are also among the brighter GCs in our sample (see Fig. 4), we believe that this is reflecting a slight brightness-size relation in this small subsample (Harris et al., 2010). The larger sample of GCs with dispersion measurements likely shows a similar size distribution as the calibration sample, but more size measurements would be needed to confirm this.
4.5 Mass-to-light ratios
The GCVD distance indicator relies on the assumption that all GCs follow a similar relation between observed light and their mass. To test this assumption, we derived dynamical mass-to-light ratios as described the -band mass-to-light ratios () using the aperture-corrected global velocity dispersions () and half-light radii () from Harris et al. (2010). We obtain virial masses for the Sombrero GCs from Eq. 1 (e.g. Larsen et al. 2002; Beasley et al. 2025):
Beasley et al. (2025) showed that Eq. 1 yields good agreement with values determined for Milky Way GCs using spatially resolved stellar dispersion profiles. Therefore, determined via Eq. 1 (e.g. also for the M31 GCs from Strader et al. 2011 and Cen A data from Dumont et al. 2022) and measured for the Milky Way GCs should be directly comparable.
The derived mass-to-light ratios are shown in Fig. 6 compared to GCs around Centaurus A (Dumont et al., 2022), M31 (Strader et al., 2011), and the Milky Way (Baumgardt et al., 2020). We show here both the GCs for which size measurements are available (larger circles) and we also included GCs that lack direct size measurements (smaller circles). For those, we assumed a size of 2.44 pc, corresponding to the median of size measurements as reported in Harris et al. (2010). These estimated values show a similar range, with several GCs having similar values as Milky Way GCs. However, we note that the uncertainties increase substantially for the fainter GCs. For M104 GCs, we find values in a range from , with a mean and a dispersion around the mean of . Many M104 GCs have higher mass-to-light ratios than the MW population, but they are also considerably brighter than most of the Galactic GCs. The brightest M104 GCs are even significantly brighter than Centauri, the most massive GC of the Milky Way, that also shows an increased mass-to-light ratio.
We find reasonable agreement with GCs in Centaurus A (), but a direct comparison is difficult based on the low number of size measurements available for our M104 sample. Similar to the behaviour seen in the Centaurus A GCs (Dumont et al., 2022), we also observe a range in mass-to-light ratio that exceeds those of the Milky Way GCs. From their analysis, Dumont et al. (2022) interpreted this scatter with a population of stripped nuclei masquerading as GCs. In M104, we find that SUCD1 indeed shows an elevated mass-to-light ratio, in line with other UCDs (e.g. Rejkuba et al. 2007; Mieske et al. 2008). Nonetheless, the defining feature that sets it apart from the GC population is its significantly larger luminosity, being about 1.4 magnitudes brighter than the next brightest GC. Based on its magnitude, size, and velocity dispersion, we derive a dynamical mass of , similar to the value reported by Hau et al. (2009).
The second panel in the bottom row of Fig. 4 shows the - plane colour-coded by the -band mass-to-light ratios. GCs with larger mass-to-light ratios appear to lie preferentially to the right of the GCVD relation. Given that the dynamical mass scales with the square of the velocity dispersion, this is perhaps not surprising, but indicates that the GCVD distance method might only by applicable for objects within a narrow range of mass-to-light ratios. Interestingly, SUCD1 is an outlier in this respect as it lies above the relation, but still has a larger mass-to-light ratio than most of the GCs. This is likely a consequence of the elevated mass-to-light ratio of SUCD1 stemming from its large size, while the higher mass-to-light ratios of the GCs are driven by an increased velocity dispersion.
4.6 Relaxation times
The final panel in Fig. 4 shows the scatter in relation to the half-mass relaxation times , estimated from the sizes and dynamical masses of the GCs. To obtain a measure of , we followed Binney and Tremaine (1987):
where ln() is used to describe the Coulomb logarithm, with as the number of stars within the GC, the average mass of a star, the mass of the GC, and the half-light radius. We assume (Giersz and Heggie, 1994), and . We find relaxation times ranging from 0.6 to 10 Gyr, while the relaxation time for SUCD1 is estimated at Gyr due to its large mass.
As the relaxation time depends on both the mass and size of the GCs, it is not surprising that the scatter in the GCVD relation shows a similar trend as with the half-light radii: smaller GCs have lower relaxation times and lie to the right of the relation.
5 Distance to Sombrero
| Photometry | Distance stat. error | No. GCs |
| Combined -band | 9.00 0.29 | 77 |
| Spitler et al. (2006) | 9.33 0.32 | 29 |
| Harris et al. (2010) | 9.00 0.33 | 29 |
| Rhode and Zepf (2004) | 8.89 0.30 | 68 |
| predicted from (Kang et al., 2022) | 9.03 0.30 | 85 |
| predicted from (Kang et al., 2022) | 9.14 0.29 | 85 |
| Source | Uncertainty | |
| Aperture/positioning error | 0.03 () | |
| Different photometric samples | 0.16 () | |
| Different | 0.20 () |
We obtained a distance to Sombrero using the GCVD approach, which is based on the fact that GCs in the Local Group obey a tight relation and on the assumption that extragalactic GC systems also follow this relation. See Beasley et al. (2024) and Beasley et al. (2025) for full details of the technique and systematic uncertainties.
5.1 Distance measurement with the GCVD relation
We used the same approach as that used in Beasley et al. (2024, 2025) and fit the relation between apparent -band magnitudes and velocity dispersions with a log-linear relation of the form:
where is the slope of the relation and describes the offset. In Beasley et al. (2024), the relation was calibrated based on Milky Way and M31 GCs, giving values of and . Therefore, when using a fixed slope, the derived offset directly scales with the distance modulus :
As described in Beasley et al. (2025), we fit the apparent -band magnitudes and velocity dispersions using a Markov Chain Monte Carlo (MCMC) ensemble sampler (emcee; Foreman-Mackey et al. 2013) to determine and its uncertainties. We used a Gaussian likelihood function that includes uncertainties in and . For the offset, we chose a uniform prior in the range of (corresponding to distances from the Milky Way out to 100 Mpc). 25 walkers and 10000 chains were used and the corresponding distances describe the median of the posterior distribution with the 16th and 84th percentiles as lower and upper uncertainties. Due to the limited range in magnitudes, we kept the slope fixed to the calibrated relation. In future work, we aim to test the universality of the slope with a larger sample of GC velocity dispersions.
We excluded SUCD1 from all distance measurements due to its peculiar properties. Additionally, it is still debated if UCDs build the high-mass end of the genuine GC population (e.g. Bekki et al. 2002; Mieske et al. 2002), or are rather the remnant nuclear star clusters (NSCs) of disrupted dwarf galaxies (e.g. Bekki et al. 2003; Pfeffer and Baumgardt 2013). While they likely are a mixed population (e.g. Norris et al. 2014; Wang et al. 2023), with its large size and high mass, SUCD1 is likely a NSC-type UCD (Hau et al., 2009). The stripping process (Pfeffer et al., 2016) as well as the complex stellar populations found in many NSCs (e.g. Kacharov et al. 2018; Fahrion et al. 2021, 2024) might explain its offset from the GCVD relation.
5.2 Distances from different samples
Distance results based on differing photometric samples are presented in Table 1. We can explore the reliability of the GCVD method by dissecting the full GC sample. For example, we can limit the GC sample to only those that have high quality HST-based Johnson -band magnitudes available. Both Spitler et al. (2006) and Harris et al. (2010) provide -band magnitudes based on the same HST/ACS imaging that comprises a 6-pointing mosaic of Sombrero. Overall the photometry from the two studies is in reasonable agreement, although there are differences in the final magnitudes due to different approaches in the aperture correction. Spitler et al. (2006) used a fixed value for all sources, while Harris et al. (2010) used size-dependent aperture corrections. While we have no specific preference for the photometry from either source, the photometry from Harris et al. (2010) might be more reliable as it takes the GC sizes into account and is based on a more detailed consideration of the point spread function that was determined empirically with proximal stars on a cluster-by-cluster basis. There are 29 GCs (excluding SUCD1) in our spectroscopic sample that overlap with both these data sets. For the photometry from Spitler et al. (2006), we measure Mpc, while for the Harris et al. (2010) photometry, we obtain Mpc.
We also matched -band photometry from the ground-based study of Rhode and Zepf (2004) with our sample, yielding 68 GCs. Rhode and Zepf (2004) adopted a single value for the extinction correction of , while Spitler et al. (2006) adopted . We therefore adjusted the Rhode and Zepf (2004) -band magnitudes by -0.008 mag to match the Spitler et al. (2006) and Harris et al. (2010) photometry.
For this sample we obtain Mpc. Therefore, all three samples give similar results, with a standard deviation of only 0.16 Mpc. In view of this, we combined the three samples taking the average of the -band magnitudes in the cases where more than one measurement was available. For this combined sample of 77 GCs, we find Mpc, which we adopt as our fiducial distance measurement.
Similarly, we can expand the sample by taking the GCs into account that have no -band magnitudes available, but photometry from Kang et al. (2022). To convert their - and -band magnitudes to -band, we made use of the EMILES SSP models (Vazdekis et al., 2010, 2016). Applying the appropriate filter transmission curves for CFHT/MegaCam and Johnson using synphot222https://synphot.readthedocs.io/en/latest/ (STScI Development Team, 2018), we derived predictions for colours depending on the metallicity of the SSP models, assuming an age of 13 Gyr representative for typical GC systems (Beasley et al., 2008; Caldwell et al., 2011). Then, we used these predictions to convert the - and -band magnitudes to -band, based on the measured metallicities (Fahrion et al., 2025). Deriving the GCVD distance with the -band magnitudes predicted from -band magnitudes yields a distance of Mpc, while applying this transformation from - to -band gives a distance of Mpc, verifying the validity of the colour conversion.
5.3 Systematics
We explored the impact of systematic effects on our distances by considering three key sources of systematic uncertainty: the impact of aperture corrections and fibre centring errors, the effect of using photometry from different sources and samples, and the impact of varying . A summary is found in Table 2.
The effect of aperture and fibre centring on the velocity dispersion measurement is discussed in Sect. 3.2. These corrections translate to an uncertainty on our distances of 0.03 Mpc (0.2%) (i.e., this is the difference in distance we obtain with GCVD between the measured velocity dispersions and the aperture-corrected dispersions). Similarly, the impact of different photometry and photometric samples is discussed in Sect. 5.2, from which we obtain a standard deviation on the three different distances of 0.16 Mpc (1.7% distance uncertainty).
To investigate the impact of varying on GCVD, we used the sizes from Harris et al. (2010) where available. To also include the remaining sample, we assumed for the other GCs a mean size of 2.44 pc, the mean of the size measurements. We then performed bootstrap resampling with replacement, where we only selected GCs with to match the Milky Way and M31 distributions. In all cases the GCVD slope was fixed to the fiducial value, and we just consider the impact on the offset . As a result of this exercise, we find that the offset can shift by up to 0.2 Mpc (2.2%).
Combining the above mentioned sources of systematics corresponds to a total systematic uncertainty of Mpc, or 2.9%.
We also explored the validity of our assumption of using a fixed slope for the GCVD relation obtained from the Milky Way and M31 data. We ran the MCMC fitting on the Sombrero data leaving both the zeropoint and the slope in Eq. 3 free. As in Beasley et al. (2024) we chose uniform priors in the range (zeropoint), and (slope). For the full sample, we found . This would translate to a distance, D = Mpc. This slope is steeper than that of the fiducial Milky Way + M31 sample () at the level. If we only select Sombrero GCs in the same magnitude range as the fiducial relation (i.e., mag) we find and obtain D = Mpc which is consistent with our distance measurement with fixed slope. To conclude, there is some evidence that the slope for the brighter Sombrero GCs might be steeper than in the Milky Way and M31 sample, but for the bulk of the sample the assumption of a fixed slope based on fiducial relation is likely reasonable and larger samples across broad magnitude ranges are required to test the effects of the slope properly.
5.4 Dependence on GC parameters
After fitting the GCVD relation to the M104 GCs, we find a scatter of the residuals of 6%. This is very similar to the scatter in the GC samples of the Milky Way, M31, and Centaurus A (see Fig. 7) and suggests that this scatter is intrinsic to such GC samples. In Fig. 8, we show the orthogonal residuals from the GCVD fit as a function of different GC parameters (similar to Fig. 4. The orthogonal residuals refer to distance in and to be best-fit line, normalised to the best-fit line.
As can be seen in this figure, the residuals show no dependence on most of the GC parameters such as metallicity, colour, line-of-sight velocity, projected radius or mass-to-light ratio. However, as also noted in Fig. 4, there is a weak relation with the size and consequently the relaxation time.
6 Comparison to literature distances
As a well studied galaxy, multiple distance measurements to M104 exist with varying accuracy. As described in Sect. 1 and discussed by McQuinn et al. (2016), distances to M104 vary between 7 and 22 Mpc, with the more accurate measurements placing the galaxy between 8 and 10 Mpc (e.g. Jensen et al. 2003; Spitler et al. 2006; McQuinn et al. 2016). McQuinn et al. (2016) note that distance measurements for M104 based on the Tully-Fisher relation generally find larger distances, between 11 – 22 Mpc (Bottinelli et al., 1984; Tully and Fisher, 1988; Sorce et al., 2014). This may be due to the peculiar morphology and spatially restricted gas distribution of M104 that make the morphological classification of this galaxy somewhat uncertain, and set it apart from more typical elliptical and spiral galaxies.
In Fig. 9, we show distances from the literature in comparison with our fiducial GCVD distance of (stat.) 0.26 (syst.) Mpc from fitting the combined -band magnitudes (see Sect. 5). We show both the statistical uncertainty and the quadrature sum of statistical and systematic uncertainty (in light grey) to visualise a total uncertainty. Distances based on the Tully-Fisher relation were excluded from Fig. 9 due to the aforementioned reasons. As can be seen from this figure, the GCVD distance compares well against literature measurements from various sources. For example, the median of the shown GCLF, PNLF, SBF, and TRGB distances is at 9.25 Mpc, and calculating the weighted mean taking the statistical uncertainties into account, we find Mpc, with a dispersion of 0.60 Mpc.
We find that the GCVD distance agrees very well with the PNLF distances reported by Ciardullo et al. (1993), Ford et al. (1996), and Ferrarese et al. (2000) that are all based on the dataset collected by Ford et al. (1996). While Ciardullo et al. (1993) used a preliminary analysis of the PNLF at that time, Ferrarese et al. (2000) re-computed the PNLF distance after deriving a Cepheid-based calibration for PNLF and other distance indicators. A similar Cepheid-based calibration was used for their SBF distance, based on HST imaging obtained with the wide field planetary camera 2 (WFPC2) described in Ajhar et al. (1997) and later used in Tonry et al. (2001). Jensen et al. (2003) used new HST Near Infrared Camera and Multi-Object Spectrometer (NICMOS) data for their SBF distance measurement, which is also in agreement with the GCVD distance.
While most distances agree with our GCVD distance within 1 of the combined statistical uncertainties, there are exceptions. For example, the GCLF distance of Mpc reported by Spitler et al. (2006) would place the Sombrero galaxy significantly closer than our GCVD distance. As Spitler et al. (2006) discuss, assuming a distance of 9.0 Mpc would make the turnover magnitude of the GCLF about 0.3 mag brighter than the universally adopted value. However, investigation of early-type galaxies in the Virgo and Fornax galaxy clusters have found that massive galaxies often have brighter GCLF turnover magnitudes (by 0.1 - 0.3 mag compared to low-mass galaxies, Villegas et al. 2010). The bright turnover magnitude in M104 might therefore be a consequence of M104 being a massive galaxy (; Jardel et al. 2011; Karunakaran et al. 2020).
Comparing the GCVD relation to the TRGB distance of Mpc from McQuinn et al. (2016) also shows an offset within 2, mostly due to the small statistical uncertainties of the TRGB measurement. Their precise value is based on dedicated, deep HST/ACS data, but McQuinn et al. (2016) quote a systematic uncertainty of Mpc stemming from the TRGB calibration (Carretta et al., 2000; Rizzi et al., 2007). Under consideration of this systematic uncertainty, the TRGB and GCVD distances agree. Sabbi et al. (2018) also derived a TRGB distance to M104 based on HST Wide Field Camera 3 (WFC3) data and they found a larger distance of Mpc. However, given that their data is shallower, they caution that the location of the TRGB is not clearly detected in their data. In the future, additional observations with JWST could be used to measure precise TRGB and SBF distances (e.g. Anand et al. 2024; Jensen et al. 2025).
7 Conclusions
In this paper, we employed the relationship between GC luminosities and their internal velocity dispersions to derive the distance to M104 using the GCVD relation. We summarise our finding as follows:
Using FLAMES high resolution spectroscopy, we derived velocity dispersions of 93 GCs and one UCD. Of those, 85 GCs have km s-1 and velocity dispersion uncertainties 5 km s-1, which we used to derive the GCVD-based distance.
Using -band magnitudes from space- and ground-based observations of 77 GCs, we applied the GCVD relation with fixed slope and obtained a distance of (stat.) 0.26 (sys.) Mpc. We find an intrinsic scatter of 6%. Testing different photometric samples including -band magnitudes derived from other filters finds consistent distance measurements.
We find that SUCD1, a known UCD near M104, does not follow the GCVD relation and instead shows an offset. This may be related to its larger size or possibly more complex stellar population than the GCs we studied. With a reported half-light radius of 14 pc, SUCD1 is significantly larger than typical GCs.
For a subsample of 29 GCs with available size measurements, we derived dynamical mass-to-light ratios, finding a range of values from . Similar values were found for GCs in the halo of Centaurus A (Dumont et al., 2022), while GCs in the Milky Way typically have lower values, but also have lower masses than the studied M104 GCs. Besides SUCD1, other GCs with elevated mass-to-light ratios follow the GCVD relation.
Our best-fitting distance of Mpc agrees well with previous distance measurements in the literature that typically range from 8 to 10 Mpc. Our GCVD distance places M104 slightly closer than the TRGB distance of Mpc reported in McQuinn et al. (2016), which currently has the lowest reported statistical uncertainty but a larger systematic uncertainty. Nonetheless, the two distance measurements agree within 2.
With velocity dispersion measurements of 85 GCs, the sample presented here is already the second largest sample of GC velocity dispersion measurements beyond the Local Group after Centaurus A. In the future, we plan to collect additional high-resolution spectroscopy of extragalactic GC systems to explore the GCVD method, its applicability and underlying systematics with larger samples.
Acknowledgments
We thank the two referees for constructive reports that have helped to improve this manuscript. KF thanks Francisco Aros for helpful discussions. KF acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101103830. ACS acknowledges support from FAPERGS (grants 23/2551-0001832-2 and 24/2551-0001548-5), CNPq (grants 314301/2021-6, 312940/2025-4, 445231/2024-6, and 404233/2024-4), and CAPES (grant 88887.004427/2024-00). A.F.M acknowledges support from RYC2021-031099-I of MICIN/AEI/10.13039/501100011033/ UE NextGenerationEU/PRTR. O.M. is grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. KF, MB, AFM, ACS, GvV, AG acknowledge support from PID2024-16088NB-I00. This work made use of Astropy:333http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013, 2018, 2022). Based on observations collected at the European Southern Observatory under ESO programme 114.274W (PI: Fahrion).
Appendix A Radial trends of magnitudes and dispersions
Figure 10 shows the -band magnitudes and velocity dispersions of the M104 GCs as a function of projected distance from the centre, illustrating that some of the brightest GCs in the sample are also among the brightest. The close match in the versus trend stems from the GCVD relation and we do not see any evidence for inflated velocity dispersions due to contamination from the galaxy light in the fiber apertures.
Appendix B Table of dispersion measurements
We report the velocity dispersions, magnitudes, and where available sizes, mass-to-light ratios, and metallicities in Table 3. The identifiers refer to the catalogue presented in Fahrion et al. (2025).
| ID1 | RA | DEC | [M/H] | Metal ref2. | ||||||
| (J2000) | (J2000) | (km s-1) | (arcsec) | (mag) | (mag) | (mag) | ( | (dex) | ||
| 12 | 190.01304 | 11.66786 | 0.34 | … | 17.46 | … | 3.380.33 | 0.480.06 | FLAMES | |
| 13 | 190.03900 | 11.64231 | 0.12 | 18.91 | 18.87 | … | 3.930.21 | 0.830.05 | FLAMES | |
| 14 | 190.04279 | 11.57125 | 0.07 | 20.11 | 20.07 | 20.14 | 1.160.47 | 1.760.11 | FLAMES | |
| 15 | 190.04479 | 11.78836 | … | … | … | 21.29 | … | 0.770.17 | FLAMES | |
| 16 | 190.04667 | 11.50436 | … | … | … | 21.21 | … | 1.140.24 | FLAMES | |
| 17 | 190.04687 | 11.84981 | … | … | … | 20.50 | … | 1.490.06 | FLAMES | |
| 18 | 190.05292 | 11.40906 | … | … | … | 21.53 | … | 0.430.23 | FLAMES | |
| 19 | 190.03271 | 11.38036 | … | … | … | 21.04 | … | 0.290.25 | FLAMES | |
| 20 | 190.05500 | 11.28775 | … | … | … | … | … | 0.510.16 | FLAMES | |
| 21 | 190.07171 | 11.95222 | … | … | … | … | … | 1.440.07 | FLAMES | |
| 23 | 190.09050 | 11.81750 | … | … | … | … | … | 0.790.17 | FLAMES | |
| 24 | 190.09571 | 11.79056 | … | … | … | 20.93 | … | 1.790.17 | FLAMES | |
| 25 | 190.13242 | 11.93833 | … | … | … | … | … | 1.750.08 | FLAMES | |
| 26 | 190.06908 | 11.55500 | … | … | … | 21.18 | … | 0.510.32 | FLAMES | |
| 27 | 190.03042 | 11.43528 | … | … | … | 20.88 | … | 2.150.15 | FLAMES | |
| 28 | 190.02520 | 11.59367 | 0.06 | 20.44 | 20.53 | 20.55 | 3.140.71 | 0.940.08 | OSIRIS | |
| 29 | 189.93025 | 11.81147 | … | … | … | 20.85 | … | 1.310.08 | FLAMES | |
| 30 | 189.95262 | 11.59462 | 0.10 | 19.09 | 19.04 | … | 2.710.17 | 0.820.09 | OSIRIS | |
| 33 | 189.97587 | 11.57515 | 0.11 | 19.93 | 19.80 | 19.83 | 3.600.57 | 1.080.12 | OSIRIS | |
| 34 | 189.97846 | 11.68622 | … | … | … | … | … | 1.140.41 | FLAMES | |
| 35 | 189.98367 | 11.71906 | … | … | … | 20.98 | … | 1.480.25 | FLAMES | |
| 36 | 190.02600 | 11.85794 | … | … | … | 20.57 | … | 1.690.14 | FLAMES | |
| 37 | 189.98450 | 11.53192 | … | … | … | … | … | 0.000.08 | FLAMES | |
| 39 | 190.02096 | 11.48725 | … | … | … | 21.00 | … | 0.400.07 | FLAMES | |
| 40 | 190.02262 | 11.93269 | … | … | … | … | … | 1.790.03 | FLAMES | |
| 41 | 190.02500 | 11.49014 | … | … | … | 21.35 | … | 1.490.17 | FLAMES | |
| 42 | 189.98825 | 11.92911 | … | … | … | … | … | 1.330.09 | FLAMES | |
| 75 | 189.98917 | 11.64292 | … | … | … | … | … | 1.850.20 | MUSE | |
| 91 | 189.99075 | 11.60258 | 0.07 | 19.39 | 19.36 | … | 1.770.21 | 1.360.13 | MUSE | |
| 134 | 190.16358 | 11.75308 | … | … | … | 20.70 | … | 1.400.13 | FLAMES | |
| 135 | 190.16233 | 11.91731 | … | … | … | 19.68 | … | 1.070.13 | FLAMES | |
| 136 | 190.15450 | 11.76708 | … | … | … | 19.29 | … | 0.980.09 | FLAMES | |
| 137 | 190.15421 | 11.49092 | … | … | … | 19.44 | … | 1.060.15 | FLAMES | |
| 139 | 190.15133 | 11.68289 | … | … | … | 20.37 | … | 1.260.05 | FLAMES | |
| 143 | 190.14529 | 11.47961 | … | … | … | 20.02 | … | 1.060.14 | FLAMES | |
| 145 | 190.14083 | 11.75581 | … | … | … | 20.45 | … | 0.930.12 | FLAMES | |
| 154 | 190.11308 | 11.47386 | … | … | … | 20.68 | … | 0.710.16 | FLAMES | |
| 160 | 190.10167 | 11.89797 | … | … | … | 21.06 | … | 1.490.03 | FLAMES | |
| 168 | 190.07933 | 11.39892 | … | … | … | 20.25 | … | 0.920.12 | FLAMES | |
| 172 | 190.07371 | 11.67914 | … | … | … | 20.23 | … | 1.260.00 | FLAMES | |
| 181 | 190.06475 | 11.60589 | 0.08 | 20.21 | 20.18 | 20.19 | 0.610.33 | 1.120.13 | FLAMES | |
| 182 | 190.06408 | 11.62633 | 0.11 | 20.76 | 20.79 | 20.62 | 2.950.80 | 1.260.00 | FLAMES | |
| 184 | 190.06367 | 11.67028 | 0.08 | 19.37 | 19.38 | 19.33 | 3.580.23 | 0.910.05 | FLAMES | |
| 189 | 190.06083 | 11.53214 | … | … | … | 19.48 | … | 0.470.15 | FLAMES | |
| 192 | 190.05729 | 11.61739 | 0.04 | 19.80 | 19.89 | 19.91 | 2.080.34 | 0.690.14 | FLAMES | |
| 194 | 190.05654 | 11.65957 | 0.07 | 19.77 | 19.77 | 19.79 | 1.760.32 | 1.120.23 | OSIRIS | |
| 198 | 190.05179 | 11.42597 | … | … | … | 19.55 | … | 0.540.12 | FLAMES | |
| 210 | 190.04362 | 11.75575 | … | … | … | 20.31 | … | 0.430.17 | FLAMES | |
| 211 | 190.04333 | 11.61078 | 0.08 | 19.02 | 18.99 | 18.98 | 1.790.14 | 0.720.07 | FLAMES | |
| 215 | 190.04037 | 11.60633 | 0.09 | 19.57 | 19.43 | 19.45 | 1.860.20 | 1.280.09 | FLAMES | |
| 218 | 190.03883 | 11.58786 | 0.06 | 19.75 | 19.83 | 19.85 | 3.510.43 | 0.570.10 | FLAMES | |
| 221 | 190.03587 | 11.65861 | 0.09 | 19.94 | 19.92 | 19.85 | 2.660.32 | 0.760.08 | FLAMES |
| ID | RA | DEC | [M/H] | Metal ref. | ||||||
| (J2000) | (J2000) | (km s-1) | (arcsec) | (mag) | (mag) | (mag) | ( | (dex) | ||
| 222 | 190.03537 | 11.72353 | … | … | … | 20.38 | … | 1.360.12 | FLAMES | |
| 226 | 190.03179 | 11.76808 | … | … | … | 20.02 | … | 0.900.11 | FLAMES | |
| 231 | 190.02949 | 11.60131 | 0.10 | 19.79 | 19.79 | 19.67 | 2.260.22 | 0.160.06 | OSIRIS | |
| 239 | 190.02342 | 11.54158 | … | … | … | 20.38 | … | 1.130.18 | FLAMES | |
| 240 | 190.02275 | 11.43819 | … | … | … | 20.06 | … | 0.500.21 | FLAMES | |
| 248 | 190.01504 | 11.48775 | … | … | … | 21.39 | … | 0.240.29 | FLAMES | |
| 249 | 190.01500 | 11.54497 | … | … | … | 20.48 | … | 0.100.14 | FLAMES | |
| 255 | 190.00900 | 11.50750 | … | … | … | 19.87 | … | 0.340.11 | FLAMES | |
| 264 | 190.00225 | 11.74897 | … | … | … | 19.77 | … | 0.210.12 | FLAMES | |
| 276 | 189.99562 | 11.58962 | 0.08 | 19.51 | 19.49 | 19.53 | 2.010.20 | 0.930.09 | OSIRIS | |
| 279 | 189.99167 | 11.57417 | 0.06 | 20.20 | 20.27 | 20.28 | 2.240.29 | 0.270.11 | FLAMES | |
| 281 | 189.98996 | 11.52369 | … | … | … | 19.28 | … | 0.400.01 | FLAMES | |
| 283 | 189.98700 | 11.33994 | … | … | … | 20.25 | … | 1.410.17 | FLAMES | |
| 288 | 189.98400 | 11.76928 | … | … | … | 20.31 | … | 1.490.04 | FLAMES | |
| 292 | 189.98221 | 11.65899 | 0.08 | 20.32 | 20.30 | 20.30 | 2.330.29 | 0.270.17 | OSIRIS | |
| 293 | 189.98083 | 11.55678 | … | … | … | 19.81 | … | 1.140.10 | FLAMES | |
| 302 | 189.97129 | 11.85125 | … | … | … | 20.66 | … | 1.110.15 | FLAMES | |
| 308 | 189.96717 | 11.63383 | 0.06 | 20.23 | 20.27 | 20.29 | 3.110.52 | 1.050.23 | FLAMES | |
| 320 | 189.96316 | 11.64266 | 0.11 | 19.92 | 19.96 | 19.70 | 3.230.32 | 0.240.05 | OSIRIS | |
| 339 | 189.94404 | 11.61414 | 0.10 | 19.84 | 19.82 | 19.68 | 2.810.32 | 0.050.10 | FLAMES | |
| 340 | 189.94371 | 11.58261 | … | … | … | 19.88 | … | 1.000.11 | FLAMES | |
| 346 | 189.93874 | 11.60014 | 0.13 | 19.33 | 19.28 | 19.13 | 3.670.26 | 0.160.04 | OSIRIS | |
| 355 | 189.92792 | 11.71342 | … | … | … | 20.38 | … | 1.490.06 | FLAMES | |
| 376 | 189.90033 | 11.56905 | … | … | … | 20.65 | … | 1.870.15 | OSIRIS | |
| 378 | 189.89896 | 11.79611 | … | … | … | 20.61 | … | 1.260.12 | FLAMES | |
| 381 | 189.89629 | 11.86453 | … | … | … | 19.98 | … | 1.100.11 | FLAMES | |
| 382 | 189.89621 | 11.50647 | … | … | … | 21.66 | … | 0.260.12 | FLAMES | |
| 387 | 189.88625 | 11.50875 | … | … | … | 20.18 | … | 0.690.13 | FLAMES | |
| 395 | 189.87221 | 11.49267 | … | … | … | 19.65 | … | 1.200.11 | FLAMES | |
| 412 | 190.04184 | 11.63909 | 0.08 | 18.79 | 18.74 | … | 3.260.26 | 0.820.10 | OSIRIS | |
| 413 | 190.00021 | 11.60261 | 0.09 | 18.94 | 18.83 | … | 2.200.30 | 0.880.10 | MUSE | |
| 414 | 190.06078 | 11.64111 | 0.11 | 18.98 | 18.83 | … | 2.290.20 | 0.570.10 | OSIRIS | |
| 423 | 189.99700 | 11.64114 | 0.04 | 20.20 | 20.29 | … | 1.970.40 | 0.220.08 | MUSE | |
| 424 | 189.96816 | 11.64635 | 0.09 | 20.20 | 20.13 | … | 2.290.61 | 1.010.13 | OSIRIS | |
| 429 | 189.97517 | 11.60142 | 0.07 | 20.44 | 20.48 | … | 2.820.73 | 0.360.23 | MUSE |